It takes Dan 2 hours to paddle his canoe 16 miles downstream, and 8 hours to return to the campsite. Find the speed of the canoe in still water.
Answer by mananth(16946)

It takes Dan 2 hours to paddle his canoe 16 miles downstream, and 8 hours to return to the campsite. Find the speed of the canoe in still water.
canoe rate =x
current rate = y
downstream rate = x+y.................2 hours..................16 miles
upstream rate = x-y...............8 hours...................16
t=d/r
16/(x+y)=2
simplify
x+y=8
16/(x-y)=8
x-y=2
add the two equations
2x=10
x=5 mph canoe speed